OpenStudy (anonymous):

Metal and Fan if x>0 and y>0 and x^9=4 and x^7=9/y^2, which of the following is an expression for the value of x in terms of y? A) 4/9 y B) 2/3 y C) 3/2 y^2 D)6y E) 36y^2

OpenStudy (bahrom7893):

\[x^9 = 4\]\[x = 4^{1/9}\]\[x^7 = 4^{7/9}\]

OpenStudy (bahrom7893):

man now I'm confused. Give me a few minutes to eat dinner, I'll get back to this later.

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

I'm seriously stuck lol

OpenStudy (bahrom7893):

Guys help!

OpenStudy (anonymous):

isolate x first then substitute?

OpenStudy (anonymous):

\[x > 0\] So obviously x is a number greater than zero. \[y > 0\] So obviously y is a number greater than zero. *WE KNOW THAT ALREADY, NOTHING NEW*

OpenStudy (bahrom7893):

We need to find x in terms of y.

OpenStudy (anonymous):

\[\large x^7=9/y^2\]\[\large x= 4^{1/9}\]\[\large 4^{1/9} =\frac{9}{y^2}\]\[\large 4^{1/9}y^2 = 9\]\[\large y^2 = \frac{9}{4^{1/9}}\]\[\large y=\ ?\]

OpenStudy (bahrom7893):

noo he needs y in terms of x.

OpenStudy (anonymous):

-_- lol.

OpenStudy (bahrom7893):

it's just a bunch of algebra, but idk why I'm stuck

OpenStudy (anonymous):

sooooo you're solving for x?

OpenStudy (anonymous):

Ohhh yeah you are lol duh

OpenStudy (anonymous):

So it's going to be solved the other way around....

OpenStudy (anonymous):

X^9 = 4 & X^7 = 9/Y^2 For the second equation multiply both sides by Y^2 to give: X^9 = 4; X^7 x Y^2 = 9 Set them so they're equal X^7 x Y^2 = [9/4] X^9 pull out the common factor X^7 Y^2 = [9/4] X^2 multiply by reciprocal and rearrange X^2 = [4/9] Y^2 take square root X = [2/3] Y -SOURCE: YAHOO